how do i solve:
3^(x-2) = 2,187??

Answer by richwmiller(17219)

3^y=2187
y=7
x-2=7
x=9
but we want to use logarithms
take the log of base 3 of both sides.
3^(x-2)=2187
x-2=7
x=9